# Multi-Objective Reinforcement Learning for Power Grid Topology Control

### v1

@@ II-B2 Topological Deviation
$R^{D}_{t}=\begin{cases}r_{\text{default}},&\text{if }D_{t}=0,\\ r_{\text{minor}},&\text{if }D_{t}\leq d_{\text{threshold}},\\ r_{\text{major}},&\text{otherwise,}\end{cases}$ (7)
where $r_{\text{default}}$ is a positive reward indicating the maintenance of the default state, $r_{\text{low}}$ is a small penalty for minor deviations from the default configuration, $r_{\text{high}}$ is a larger penalty for significant deviations, and $d_{\text{threshold}}$ is the predefined threshold for minor deviations.
@@ II-B3 Switching Frequency
$R^{F}_{t}=\begin{cases}r_{\text{DoNothing}},&\text{if }F_{t}=0,\\ r_{\text{low}},&\text{if }F_{t}\leq F_{\text{low}}^{\text{th}},\\ r_{\text{high}},&\text{if }F_{t}\geq F_{\text{high}}^{\text{th}},\end{cases}$ (9)
where $r_{\text{default}}$ is a baseline reward indicating no switching actions, $r_{\text{low}}$ is a penalty for low switching frequency, $r_{\text{high}}$ is a penalty for high switching frequency, and $F_{\text{low}}^{\text{th}},F_{\text{high}}^{\text{th}}$ are the predefined thresholds for low and high switching frequencies, respectively.
@@ II-C Deep Optimistic Linear Support
To generate a set of policies, we employ the deep optimistic linear support (DOL) algorithm to iteratively construct the optimal solution set of policies [24]. DOL functions as an outer-loop MORL approach [25], leveraging optimistic linear support (OLS) [27] to iteratively construct the convex coverage set (CCS) of a multi-objective problem by generating scalarization functions in the form of weight vectors $\mathbf{w}$ and updating the solution set $\Omega^{s}$ accordingly [25]. Additionally, here DOL facilitates integration of MOPPO as a multi-objective single policy RL Algorithm.
Algorithm 2 provides details of the proposed DOL approach. In the first iteration, DOL assigns the extrema weights to the queue with infinite priority, ensuring these weights are processed first by the MOPPO algorithm (line 7). For each new iteration of the DOL algorithm, the weight vector with the highest priority $q_{max}$ (expected highest improvement) will be given to the MOPPO (line 10). The MOPPO is trained using $\mathbf{w}$, described in Algorithm 1. The trained agent containing $\theta_{new}$ is evaluated and produces $\mathbfcal{V}$ as the average value vector across the evaluation episodes (line 21). $\mathbfcal{V}$ is used to evaluate if the trained model contributes to the CCS. $\mathbfcal{V}$ is the sampled vectorized value-estimation $\mathbf{V}(s)$ over the evaluation episodes, based on the received rewards; hence, $\mathbfcal{V}$ does not necessarily reflect the vectorized state dependent value function $\mathbf{V}(s)$ (1).
@@ III-A Settings
All case studies are performed on the RTE 5-bus system in the Grid2Op environment [26], providing initial insights into a MORL approach for topology control. Figure 2 shows the 5-bus system. Each environment scenario lasts a week with a 5-min resolution (2016 time steps), which corresponds to the maximum episode duration ($E$). We use 16 scenarios for training, 2 scenarios for validation, and 2 scenarios for testing. The experiments are performed on up to 20 random seeds initializing the environment. The PPO neural network architecture consists of 2 fully connected layers with 64-dimensional hidden features. Training is performed using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 512. The MOPPO algorithm assumes 4 update cycles. In the case study on robustness to contingencies (Section III-C), an adversarial agent is considered to simulate N-1 contingency states by randomly targeting power lines. Additionally, a set of common expert rules are considered to improve the performance and ensure safety, as detailed in [14, 8]. All computations are performed using DelftBlue’s supercomputer, equipped with Intel XEON E5-6248R 24C 3.0GHz CPU cores [28]. We use Grid2Op 1.10, LightSim2Grid 0.8, pandapower 2.14, gymnasium 0.29, mo-gymnasium 1.1 and the morl-baselines 1.0 package. The code for this study is publicly available in [29].
@@ Acknowledgment
AI4REALNET has received funding from European Union’s Horizon Europe Research and Innovation programme under the Grant Agreement No 101119527. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ II-B2 Topological Deviation
$R^{D}_{t}=\begin{cases}{\color[rgb]{0,0,1}r_{\text{default}}^{\text{D}}},&\text{if }D_{t}=0,\\ {\color[rgb]{0,0,1}r_{\text{low}}^{\text{D}}},&\text{if }D_{t}\leq d_{\text{threshold}},\\ {\color[rgb]{0,0,1}r_{\text{high}}^{\text{D}}},&\text{otherwise,}\end{cases}$ (7)
where ${\color[rgb]{0,0,1}r_{\text{default}}^{\textbf{D}}}$ is a positive reward indicating the maintenance of the default state, ${\color[rgb]{0,0,1}r_{\text{\text{low}}}^{\text{D}}}$ is a small penalty for minor deviations from the default configuration, ${\color[rgb]{0,0,1}r_{\text{high}}^{\text{D}}}$ is a larger penalty for significant deviations, and $d_{\text{threshold}}$ is the predefined threshold for minor deviations.
@@ II-B3 Switching Frequency
$R^{F}_{t}=\begin{cases}{\color[rgb]{0,0,1}r_{\text{default}}^{\text{F}}},&\text{if }F_{t}=0,\\ {\color[rgb]{0,0,1}r_{\text{low}}^{\text{F}}},&\text{if }F_{t}\leq F_{\text{low}}^{\text{th}},\\ {\color[rgb]{0,0,1}r_{\text{high}}^{\text{F}}},&{\color[rgb]{0,0,1}\text{otherwise}}\end{cases}$ (9)
where ${\color[rgb]{0,0,1}r_{\text{default}}^{\text{F}}}$ is a baseline reward indicating no switching actions, ${\color[rgb]{0,0,1}r_{\text{low}^{\text{F}}}}$ is a penalty for low switching frequency, ${\color[rgb]{0,0,1}r_{\text{high}}^{\text{F}}}$ is a penalty for high switching frequency, and $F_{\text{low}}^{\text{th}}$ is the predefined threshold for low and high switching frequencies, respectively. $R^{F}_{t}$ focuses on the operational interactions in the grid e.g. material wear off and effort for operators, whereas $R^{D}_{t}$ is directed to operational robustness of the power grid [4].
@@ II-C Deep Optimistic Linear Support
To generate a set of policies, we employ the deep optimistic linear support (DOL) algorithm to iteratively construct the optimal solution set of policies [24]. DOL functions as an outer-loop MORL approach [25], leveraging optimistic linear support (OLS) [27] to iteratively construct the convex coverage set (CCS) of a multi-objective problem by generating scalarization functions in the form of weight vectors $\mathbf{w}$ and updating the solution set $\Omega^{s}$ accordingly [25]. This outerloop approach is choosen, as to incorporate already existing research in single objective RL algorithms which solve the topology control problem[24]. Algorithm 2 provides details of the proposed DOL approach. In the first iteration, DOL assigns the extrema weights to the queue with infinite priority, ensuring these weights are processed first by the MOPPO algorithm (line 7). For each new iteration of the DOL algorithm, the weight vector with the highest priority $q_{max}$ (expected highest improvement) will be given to the MOPPO (line 10). The MOPPO is trained using $\mathbf{w}$, described in Algorithm 1. The trained agent containing $\theta_{new}$ is evaluated and produces $\mathbfcal{V}$ as the average value vector across the evaluation episodes (line 21). $\mathbfcal{V}$ is used to evaluate if the trained model contributes to the CCS. $\mathbfcal{V}$ is the sampled vectorized value-estimation $\mathbf{V}(s)$ over the evaluation episodes, based on the received rewards; hence, $\mathbfcal{V}$ does not necessarily reflect the vectorized state dependent value function $\mathbf{V}(s)$ (1).
@@ III-A Settings
All case studies are performed on the RTE 5-bus system in the Grid2Op environment [26], providing initial insights into a MORL approach for topology control. As this is the first study on the inquiry of MORL on topology control we focus on the RTE 5-bus system. Figure 2 shows the 5-bus system. Each environment scenario lasts a week with a 5-min resolution (2016 time steps), which corresponds to the maximum episode duration ($E$). As only little scenarios are available, we use 16 scenarios for training, 2 scenarios for validation, and 2 scenarios for testing. The data used is available, when initializing the respective environment from Grid2Op [26]. The experiments are performed on up to 20 random seeds initializing the environment. The PPO neural network architecture consists of 2 fully connected layers with 64-dimensional hidden features. Training is performed using the Adam optimizer with a learning rate of $5\times 10^{-4}$ and a batch size of 512. The MOPPO algorithm assumes 4 update cycles. Initial hyperparameters are derived from [19] and optimized using empirical experience. In the case study on robustness to contingencies (Section III-C), an adversarial agent is considered to simulate N-1 contingency states by randomly targeting power lines. Additionally, a set of common expert rules are considered to improve the performance and ensure safety, as detailed in [14, 8]. All computations are performed using DelftBlue’s supercomputer, equipped with Intel XEON E5-6248R 24C 3.0GHz CPU cores [28]. We use Grid2Op 1.10, LightSim2Grid 0.8, pandapower 2.14, gymnasium 0.29, mo-gymnasium 1.1 and the morl-baselines 1.0 package. The code for this study is publicly available in [29].
@@ Acknowledgment
This work was supported through the AI-EFFECT project (Grant Agreement No 101172952) and AI4REALNET (Grant Agreement No 101119527) ); funded under the European Union’s Horizon Europe Research and Innovation program. However, views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union.